# Context versus Prior Knowledge in Language Models

### v1

@@ E.1 Relevant vs Irrelevant Context Persuasion Scores Across Models
Our null hypothesis that the mean persuasion score of relevant contexts is not greater than that of irrelevant contexts. We summarize the test results (effect size and p-values) for all queries for all models in Fig. 13.
@@ E.2 Assertive vs Base Context Persuasion Scores Across Models
Our null hypothesis that the mean persuasion score of assertive contexts is not greater than that of base contexts. We summarize the test results (effect size and p-values) for all queries for all models in Fig. 14.
@@ E.3 Negation vs Base Context Persuasion Scores Across Models
Our null hypothesis that the mean persuasion score of negation contexts is not equal to that of base contexts. We summarize the test results (effect size and p-values) for all queries for all models in Fig. 15.
@@ F.1 Unfamiliar vs Familiar Entity Susceptibility Scores Across Models
Our null hypothesis that the mean susceptibility score of unfamiliar entities is not greater than that of familiar entities. We summarize the test results (effect size and p-values) for all queries for all models in Fig. 16. From this figure, we can see the trend of how effect size and percentage of significant queries generally increases with model size, and notably the smallest model has no significant results for any query. However, even the larger models do not exhibit significant differences in scores between unfamiliar and familiar entities for all queries. To investigate the spread further, we plot the p-values and effect size against the entity frequencies in the Pile. The results are presented in Fig. 17. There is a significant trend for the open queries against the frequency (spearman, is , ), showing that real entities tend to be less susceptible the more frequently they appear in the training data. The trend is not significant for the closed set.
@@ F.2 Training Data and Susceptibility Scores
Since language models are parameterized with knowledge from their training corpora, we examine whether we can identify correlations between patterns in entity susceptibility scores and their prevalence in the training data.
Since the Pythia models are all trained on the Pile dataset Gao et al. (2020) in a single pass, we choose to compare the susceptibility scores from the Pythia models to various frequency and co-occurrence statistics in the Pile.141414We use the same deduplicated 825GiB version of the Pile that the Pythia-6.9b model was trained on. https://huggingface.co/datasets/EleutherAI/the_pile_deduplicated.
@@
Our goal is to understand how the susceptibility score relates to the frequencies of entities and their co-occurrences in the training data. For this, we use all of the entities and answers selected as described in § 4.4 and locate them in the Pile. We only perform rudimentary tokenization of each document by removing punctuation and splitting at white spaces, but find this suffices to locate the exact terms. As a sanity check, we annotate named entities in 30k documents and cross-reference the list of entities. If a supposed entity has a fairly high frequency (>50) and is most often (>75%) not labeled as an entity, we exclude it from the calculations. This removes 200 entities that are high-frequency non-entity words in English and lowers the co-occurrences by 25%. Finally, we calculate the token distance for each entity-answer pair for every document in 1/3 of the Pile.
@@ F.3 Entity Degrees and Susceptibility Scores
Since knowledge graphs are structured conceptual maps relating entities, we further seek to identify whether we can identify any correlations between statistics of the YAGO knowledge graph and an entity's susceptibility scores.
Validating against a knowledge graph can be advantageous over validating against the actual training data for a number of reasons, including that for most models, the actual training data is inaccessible for research, and the scale of the training data can make it prohibitively expensive to trawl through efficiently and precisely. For example, with a knowledge graph, we can identify the exact number of objects an entity might share an alumniOf relation with, while within the pretraining data, it is very difficult to identify the number of different answers an entity will co-occur within the context of the specific alumniOf relation.
@@ G.1 Social Sciences Measurement
Large language models (LLMs) are actively used today in empirical social sciences for annotating data and descriptive data analysis (e.g., classifying tweets with sentiment and ideology scores (Ziems et al., 2024; Gilardi et al., 2023)). However, Zhang et al. (2023) warn that LLMs applied to sentiment classification ``may inadvertently adopt human biases'' and demonstrate that the prompt design, i.e., the context persuasiveness, can significantly influence the outcome. O'Hagan and Schein (2023) demonstrate that LLMs exhibit biases about different entities when measuring political ideology, based on their prior knowledge. Finally, Stoehr et al. (2024) while use LLMs to measure the stance of product reviews, their setting does not disentangle the effects of prior knowledge and context, thus leaving ambiguous the question of whether the measurement is more because of the LLM's prior bias about the product or the review's actual content.
@@
We consider a manually constructed dataset (Tab. 1) of well-known entity-pairs which are either friends (e.g., Harry Potter and Ron Weasley) or enemies (e.g., David and Goliath) and contexts relating the pair (e.g., Harry loves Ron). We aim to understand how susceptibility scores may differ between the two kinds of relationships for the query What's the relationship between {entity1} and {entity2}?, e.g., are famous friend-based relationships more susceptible than enemy-based relationships? We compute the susceptibility scores for these entity-pairs using simple template-generated contexts such as {entity1} loves {entity2} and {entity1} hates {entity2} and then analyze the results.
@@ Results.
Results are shown in Fig. 12. For all contexts, there is a significant difference between the genders (). We make several observations: (i) When prompted with the masculine stereotypes, we get the lowest susceptibility for the male entities, indicating the model has a bias for the stereotypical male contexts. (ii) By swapping the genders in the stereotypical female contexts (M*), we get the highest susceptibility for the male entities. Somewhat surprisingly, we also find that this gives the lowest susceptibility scores for the female names. (iii) The susceptibility scores for the female names over the stereotypical female contexts (F) are similar to the swapped ones (F*), hinting that the model does not have as strong a bias for the feminine stereotypes as in the masculine case in (i). Combined with our prior findings on the relation between susceptibility and the saliency of entities, this hints at masculine stereotypes being more prominently represented in the training data (as well as in the model's induced biases). At the same time, we don't see evidence of exaggerated feminine stereotypes being favoured via lower susceptibility scores.
@@ G.2.1 Gender Stereotypes Data
Here, we list the data used to run the gender bias use-case. All data was generated with the help of the GPT-4 LLM.
@@
The stereotypical feminine contexts are: Women are not as strong as men., All women are overly emotional and sensitive., Women are not good at math and sciences., All women are naturally good at cooking and cleaning., Women talk more than men., Women are generally not capable of doing manual labor., All women love shopping and fashion., Women don't understand technology as well as men do., Women are indecisive and change their mind too often., Women cannot drive as well as men., Women are not as ambitious as men when it comes to career., Women overthink things more than men., Women are not as good at sports as men., Women are more gossipy than men., All women are afraid of bugs and spiders., Women are not good at directions and get lost easily., Women take longer to get ready than men., Women always need to ask for others for their opinion., All women are obsessed with their weight and dieting., All women are crazy about babies and want to have kids.
The stereotypical masculine contexts are Men always bring the association of strength and aggression., All men are inherently messy and cannot keep their spaces clean., Men are not good at multitasking., All men are naturally good at sports., Men cannot cook as well as women., Men do not take care of their health as much as women do., I believe men are not as talkative as women., All men love video games., Men are not good listeners., Men are always competitive., Men are not as emotional as women., Men are not able to express their feelings as well as women do., Men only think about one thing., Men look at problems from a logical perspective, not considering emotions., Men are typically less organized than women., Men tend to make decisions without considering others' feelings., All men only care about sports and cars., Men don't like to ask for help., Men don't pay as much attention to appearance as women do., All men can't resist a flirt.
The queries used for both sets of entities are: I think {entity} works as a, {entity} is good at, I met {entity} yesterday, they are, {entity} usually spends their free time, {entity}'s favorite hobby is likely, {entity} probably lives in a, When it comes to {entity}'s fashion style, they prefer, {entity}'s educational background might be in, {entity} often enjoys eating, If {entity} was a colour, they would be, The one thing that {entity} cannot resist is, {entity} probably drives a, One thing that {entity} is really bad at is, {entity}'s personality can be described as, {entity} spends most of their time, {entity} probably likes to watch, During weekends, {entity} might be, {entity}'s favorite type of music is probably, I imagine {entity} has a pet, it is a, {entity}'s favorite type of movie is probably.

### v3

@@ E.1 Relevant vs Irrelevant Context Persuasion Scores Across Models
Our null hypothesis is that the mean persuasion score of relevant contexts is not greater than that of irrelevant contexts. We summarize the test results (effect size and p-values) for all queries for all models in Fig. 12.
@@ E.2 Assertive vs Base Context Persuasion Scores Across Models
Our null hypothesis is that the mean persuasion score of assertive contexts is not greater than that of base contexts. We summarize the test results (effect size and p-values) for all queries for all models in Fig. 13.
@@ E.3 Negation vs Base Context Persuasion Scores Across Models
Our null hypothesis is that the mean persuasion score of negation contexts is not equal to that of base contexts. We summarize the test results (effect size and p-values) for all queries for all models in Fig. 14.
@@ F.1 Unfamiliar vs Familiar Entity Susceptibility Scores Across Models
Our null hypothesis is that the mean susceptibility score of unfamiliar entities is not greater than that of familiar entities. We summarize the test results (effect size and p-values) for all queries for all models in Fig. 15. From this figure, we can see the trend of how effect size and percentage of significant queries generally increase with model size, and notably the smallest model has no significant results for any query. However, even the larger models do not exhibit significant differences in scores between unfamiliar and familiar entities for all queries. To investigate the spread further, we plot the p-values and effect size against the entity frequencies in the Pile. The results are presented in Fig. 16. There is a significant trend for the open queries against the frequency (spearman, is , ), showing that real entities tend to be less susceptible the more frequently they appear in the training data. The trend is not significant for the closed set.
@@ F.2 Training Data and Susceptibility Scores
Since language models are parameterized with knowledge from their training corpora, we examine whether we can identify correlations between patterns in entity susceptibility scores and their prevalence in the training data. Because the Pythia models are all trained on the Pile dataset Gao et al. (2020) in a single pass, we choose to compare the susceptibility scores from the Pythia models to various frequency and co-occurrence statistics in the Pile.141414We use the same deduplicated 825GiB version of the Pile that the Pythia-6.9b model was trained on. https://huggingface.co/datasets/EleutherAI/the_pile_deduplicated.
@@
Our goal is to understand how the susceptibility score relates to the frequencies of entities and their co-occurrences in the training data. For this, we use all of the entities and answers selected as described in § 4.4 and locate them in the Pile. We only perform rudimentary tokenization of each document by removing punctuation and splitting at white spaces, but find this suffices to locate the exact terms. As a sanity check, we annotate named entities in 30k documents and cross-reference the list of entities. If a supposed entity has a fairly high frequency (¿50) and is most often (¿75%) not labeled as an entity, we exclude it from the calculations. This removes 200 entities that are high-frequency non-entity words in English and lowers the co-occurrences by 25%. Finally, we calculate the token distance for each entity-answer pair for every document in 1/3 of the Pile.
@@ F.3 Entity Degrees and Susceptibility Scores
Since knowledge graphs are structured conceptual maps relating entities, we further seek to identify whether we can identify any correlations between statistics of the YAGO knowledge graph and an entity’s susceptibility scores. Validating against a knowledge graph can be advantageous over validating against the actual training data for a number of reasons, including that for most models, the actual training data is inaccessible for research, and the scale of the training data can make it prohibitively expensive to trawl through efficiently and precisely. For example, with a knowledge graph, we can identify the exact number of objects an entity might share an alumniOf relation with, while within the pretraining data, it is very difficult to identify the number of different answers an entity will co-occur within the context of the specific alumniOf relation.
@@ G.1 Social Sciences Measurement
Large language models (LLMs) are actively used today in empirical social sciences for annotating data and descriptive data analysis (e.g., classifying tweets with sentiment and ideology scores (Ziems et al., 2024; Gilardi et al., 2023)). However, Zhang et al. (2023) warn that LLMs applied to sentiment classification “may inadvertently adopt human biases” and demonstrate that the prompt design, i.e., the context persuasiveness, can significantly influence the outcome. O’Hagan and Schein (2023) demonstrate that LLMs exhibit biases about different entities when measuring political ideology, based on their prior knowledge. Finally, Stoehr et al. (2024) use LLMs to measure the stance of product reviews, their setting does not disentangle the effects of prior knowledge and context, thus leaving ambiguous the question of whether the measurement is more because of the LLM’s prior bias about the product or the review’s actual content.
@@
We consider a manually constructed dataset (Tab. 1) of well-known entity-pairs which are either friends (e.g., Harry Potter and Ron Weasley) or enemies (e.g., David and Goliath) and contexts relating the pair (e.g., Harry loves Ron). We aim to understand how susceptibility scores may differ between the two kinds of relationships for the query What’s the relationship between {entity1} and {entity2}?, e.g., are famous friend-based relationships more susceptible than enemy-based relationships? We compute the susceptibility scores for these entity-pairs using simple template-generated contexts such as {entity1} loves {entity2} and {entity1} hates {entity2} and then analyze the results.
@@ Results.
Results are shown in Fig. 11. For all contexts, there is a significant difference between the genders (). We make several observations: (i) When prompted with the masculine stereotypes, we get the lowest susceptibility for the male entities, indicating the model has a bias for the stereotypical male contexts. (ii) By swapping the genders in the stereotypical female contexts (M*), we get the highest susceptibility for the male entities. Somewhat surprisingly, we also find that this gives the lowest susceptibility scores for the female names. (iii) The susceptibility scores for the female names over the stereotypical female contexts (F) are similar to the swapped ones (F*), hinting that the model does not have as strong a bias for the feminine stereotypes as in the masculine case in (i). Combined with our prior findings on the relation between susceptibility and the saliency of entities, this hints at masculine stereotypes being more prominently represented in the training data (as well as in the model’s induced biases). At the same time, we don’t see evidence of exaggerated feminine stereotypes being favored via lower susceptibility scores.
@@ G.2.1 Gender Stereotypes Data
Here, we list the data used to run the gender bias use case. All data was generated with the help of the GPT-4 LLM.
@@
The stereotypical feminine contexts are: Women are not as strong as men., All women are overly emotional and sensitive., Women are not good at math and sciences., All women are naturally good at cooking and cleaning., Women talk more than men., Women are generally not capable of doing manual labor., All women love shopping and fashion., Women don’t understand technology as well as men do., Women are indecisive and change their mind too often., Women cannot drive as well as men., Women are not as ambitious as men when it comes to career., Women overthink things more than men., Women are not as good at sports as men., Women are more gossipy than men., All women are afraid of bugs and spiders., Women are not good at directions and get lost easily., Women take longer to get ready than men., Women always need to ask for others for their opinion., All women are obsessed with their weight and dieting., All women are crazy about babies and want to have kids.
The stereotypical masculine contexts are Men always bring the association of strength and aggression., All men are inherently messy and cannot keep their spaces clean., Men are not good at multitasking., All men are naturally good at sports., Men cannot cook as well as women., Men do not take care of their health as much as women do., I believe men are not as talkative as women., All men love video games., Men are not good listeners., Men are always competitive., Men are not as emotional as women., Men are not able to express their feelings as well as women do., Men only think about one thing., Men look at problems from a logical perspective, not considering emotions., Men are typically less organized than women., Men tend to make decisions without considering others’ feelings., All men only care about sports and cars., Men don’t like to ask for help., Men don’t pay as much attention to appearance as women do., All men can’t resist a flirt.
The queries used for both sets of entities are: I think {entity} works as a, {entity} is good at, I met {entity} yesterday, they are, {entity} usually spends their free time, {entity}’s favorite hobby is likely, {entity} probably lives in a, When it comes to {entity}’s fashion style, they prefer, {entity}’s educational background might be in, {entity} often enjoys eating, If {entity} was a colour, they would be, The one thing that {entity} cannot resist is, {entity} probably drives a, One thing that {entity} is really bad at is, {entity}’s personality can be described as, {entity} spends most of their time, {entity} probably likes to watch, During weekends, {entity} might be, {entity}’s favorite type of music is probably, I imagine {entity} has a pet, it is a, {entity}’s favorite type of movie is probably.